# Maxime Dieudonné1, Guillaume Auzias1,†, Julien Lefèvre1,†

### v1

@@ 5.1 The importance of global normalization
The neuroimaging literature is increasingly invoking the sulcal depth as a relevant feature for characterizing brain morphology, in particular during development. Sulcal depth has been used to describe the spatio-temporal evolution of brain folding in human fetuses and newborns [Yun et al., b, Xu et al.,, Lefèvre et al., c], and between birth and adulthood [Hill et al.,]. Recent studies suggest that sulcal depth in fetuses provides prognostic value of later cognitive development [Bartha-Doering et al.,, Hahner et al.,], and is instrumental to detect neuro-developmental malformations [Im et al., b, Brun et al.,]. Sulcal depth is also used to validate biomechanical models of gyrification [Wang et al.,, Alenyá et al.,]. In all these exemplar studies, the use of our $DPF^{*}$ would enable better interpretation.
Indeed, when confronted to the issues related to comparing morphometric features between brains of different sizes, various strategies have been experimented. For instance, in [Yun et al., b], the authors compared cortical folding patterns across fetuses at different ages, and with a group of adults. Smoothed mean curvature was used as a proxy of sulcal depth in fetal brains in order to detect small and shallow folds, while the method in [Yun et al., a] was used in adults. The use of different methods at different ages limits the interpretation to large effects. In [Meng et al.,], the authors decomposed the cortex in sulcal basins. To compensate for variations in brain size, the parameters of their algorithm were adapted using linear regression models including as main factors surface area and maximum depth. In these works, the normalisation is empirical, and the improvements compared to non-normalised estimation techniques is only qualitative. The $DPF^{*}$ proposed in this work enables to leverage these serious limitations.
@@ 5.2 Literature on formal normalisation of surface features
To our knowledge, the notion of scale invariance as a ”desirable property of measures of brain shape” was first proposed in [Batchelor et al.,]. The authors illustrated the problem for curvature and various gyrification indices. Similarly in [Pienaar et al.,] the authors documented the changes in growth and folding during the development, and their impact on curvature measures. However these seminal works did not propose any implementable solution. A pragmatic and formal strategy called Nondimensionalisation was proposed in [Rodriguez-Carranza et al.,] and used in [Knutsen et al.,]. The principle consists in dividing a folding measure by an estimate of global brain size in order to obtain a dimensionless quantity. This global metric was denoted as normalisation factor and defined as $3V/A$ in [Rodriguez-Carranza et al.,] while in [Knutsen et al.,] it was called characteristic length and defined as $\sqrt{A/4\pi}$. These two formula are equivalent in the case of a perfect sphere and correspond to its radius. In the case of real brains, however, normalising with respect to either the volume or the area of the surface will not be equivalent since surface area and volume are related by an allometric rule, as pointed in the introduction. In the absence of better understanding of the causal mechanisms relating brain maturation and folding, normalising using either surface are or volume amounts to making an assumption in favour of one theoretical model of brain folding against the others [Ronan and Fletcher,]. Our method also requires the estimation of a normalisation factor, we are thus confronted to the same question. In this work, we used the cubic root of the volume. This choice is based on the assumption that brain folding is a consequence of a mechanical instability due to brain growth [Tallinen et al.,], which implies that the increase of the surface area is a consequence of the increase in tissues volume.
To our knowledge, our study is the first to extend the normalisation of brain measures beyond those defined by simple formula, and in particular to partial differential equations. The nondimensionalisation approach we used is relatively standard in the field of differential equations coming from the physics [Sánchez Pérez et al.,]. It implies considering characteristic quantities of the problem, e.g. with respect to space or time. In our case we only needed to define a characteristic length. An interesting extension of our approach could be to incorporate a characteristic time involved in a spatio-temporal evolution, such as surface smoothing as proposed in [Lefèvre et al., b].
@@ 5.3 Tuning of the parameter for $DPF^{*}$
Like the original $DPF$ method in [Boucher et al.,], the $DPF^{*}$ depends on a parameter $\alpha$ which has the physical unit of the inverse of an area ($m^{-2}$). We showed in section 2.3 the relationship between this parameter and the size of a spatial kernel: a larger parameter value induces smaller spatial neighbourhood influencing the measure and vice versa. In our approach, the scale-control property enables to obtain a similar regularisation level despite variations in brain size. To our opinion it is a crucial point which deserves more attention in neuroimaging where smoothing procedures are rarely used with an adaptive kernels for various brain sizes, except for instance [Kim et al.,].
5.4 $DPF*$ does not compensate for variations beyond global scaling
Our approach is the first to explicitly allow for compensating for the influence of global brain size. However, the increasing complexity of cortical folding during development [Lefèvre et al., a] or allometric variations with respect to brain size between adults [Toro et al.,, Reardon et al.,] suggest that more subtle empirical laws exist between sulcal depth and the global size of the brain. Specifically, our observation of a strong influence of brain size (higher Wasserstein distance) on the $DPF*$ for the 200 smaller, younger subjects despite our formal normalization is expected. Indeed, during early brain development, the increase in size is accompanied by a major increase in folding. Normalizing for global scaling should thus not fully compensate for variations in sulcal depth. The ultimate sulcal depth estimation technique would eventually consider local normalization factors. Designing such a method allows us to also compensate for local variations in cortical geometry, for instance using the approach introduced in [Joshi et al.,]. However, note that modelling variations in local cortical folding would amount to propose a model of brain folding like for instance [Tallinen et al.,]. The inclusion of local normalization factors thus corresponds to a major change in the objectives and perspectives of the approach.
@@ 6.6 Qualitative comparison on a spatio-temporal atlas
In order to complement the quantitative results and observations from the previous experiments, we applied the different estimation techniques to the spatio-temporal surfacic atlas computed from the dHCP dataset in [Bozek et al.,]. This atlas is representative of the evolution of the cortical surface between 28 and 44 weeks post-conception and thus shows marked changes in both the global size and in the magnitude of folding.
We show in Fig. 8 the sulcal depth maps obtained using $DPF^{*}$ and SULC on the surfacic atlas from the dHCP [Bozek et al.,] at 28, 32 and 44 weeks post conception. We observe a better dissociation between sulci and gyri at every age for $DPF^{*}$ compared to SULC. In addition, the estimated depth values are more homogeneous gyral crests distributed across cortical regions with $DPF^{*}$, while with SULC we observe that high depth values are concentrated around large convex regions such as the anterior temporal lobe, the frontal pole and the occipital region. The $DPF^{*}$ gives a finer description of the folds, with a better definition of the small sulci located in the inferior temporal region, as well as in the medial side of the cortex (see for instance the small folds above the cingular sulcus).
Figure 8: Application of $DPF^{*}$ (left) and SULC (right) to the spatio-temporal atlas from [Bozek et al.,], computed from the dHCP dataset corresponding to 28, 32 and 44 weeks post-conception from top to bottom.

### v2

@@ 5.1 The importance of global normalization
The neuroimaging literature is increasingly invoking the sulcal depth as a relevant feature for characterizing brain morphology, in particular during development. Sulcal depth has been used to describe the spatio-temporal evolution of brain folding in human fetuses and newborns yun_temporal_2020; xu_spatiotemporal_2022; lefevre_identification_2009, and between birth and adulthood hill_similar_2010. Recent studies suggest that sulcal depth in fetuses provides prognostic value of later cognitive development bartha-doering_fetal_2023; hahner_global_2019, and is instrumental to detect neuro-developmental malformations im_atypical_2016; brun_localized_2016. Sulcal depth is also used to validate biomechanical models of gyrification wang_influence_2021; alenya_computational_2022. In all these exemplar studies, the use of our $DPF^{*}$ would enable better interpretation.
Indeed, when confronted to the issues related to comparing morphometric features between brains of different sizes, various strategies have been experimented. For instance, in yun_temporal_2020, the authors compared cortical folding patterns across fetuses at different ages, and with a group of adults. Smoothed mean curvature was used as a proxy of sulcal depth in fetal brains in order to detect small and shallow folds, while the method in yun_automated_2013 was used in adults. The use of different methods at different ages limits the interpretation to large effects. In meng_spatial_2014, the authors decomposed the cortex in sulcal basins. To compensate for variations in brain size, the parameters of their algorithm were adapted using linear regression models including as main factors surface area and maximum depth. In these works, the normalisation is empirical, and the improvements compared to non-normalised estimation techniques is only qualitative. The $DPF^{*}$ proposed in this work enables to leverage these serious limitations.
@@ 5.2 Literature on formal normalisation of surface features
To our knowledge, the notion of scale invariance as a ”desirable property of measures of brain shape” was first proposed in batchelor_measures_2002. The authors illustrated the problem for curvature and various gyrification indices. Similarly in pienaar_methodology_2008 the authors documented the changes in growth and folding during the development, and their impact on curvature measures. However these seminal works did not propose any implementable solution. A pragmatic and formal strategy called Nondimensionalisation was proposed in rodriguez-carranza_framework_2008 and used in knutsen_spatial_2013. The principle consists in dividing a folding measure by an estimate of global brain size in order to obtain a dimensionless quantity. This global metric was denoted as normalisation factor and defined as $3V/A$ in rodriguez-carranza_framework_2008 while in knutsen_spatial_2013 it was called characteristic length and defined as $\sqrt{A/4\pi}$. These two formula are equivalent in the case of a perfect sphere and correspond to its radius. In the case of real brains, however, normalising with respect to either the volume or the area of the surface will not be equivalent since surface area and volume are related by an allometric rule, as pointed in the introduction. In the absence of better understanding of the causal mechanisms relating brain maturation and folding, normalising using either surface are or volume amounts to making an assumption in favour of one theoretical model of brain folding against the others ronan_genes_2015. Our method also requires the estimation of a normalisation factor, we are thus confronted to the same question. In this work, we used the cubic root of the volume. This choice is based on the assumption that brain folding is a consequence of a mechanical instability due to brain growth tallinen_growth_2016, which implies that the increase of the surface area is a consequence of the increase in tissues volume.
To our knowledge, our study is the first to extend the normalisation of brain measures beyond those defined by simple formula, and in particular to partial differential equations. The nondimensionalisation approach we used is relatively standard in the field of differential equations coming from the physics sanchez_perez_searching_2017. It implies considering characteristic quantities of the problem, e.g. with respect to space or time. In our case we only needed to define a characteristic length. An interesting extension of our approach could be to incorporate a characteristic time involved in a spatio-temporal evolution, such as surface smoothing as proposed in lefevre_surface_2013.
@@ 5.3 Tuning of the parameter for $DPF^{*}$
Like the original $DPF$ method in boucher_depth_2009, the $DPF^{*}$ depends on a parameter $\alpha$ which has the physical unit of the inverse of an area ($m^{-2}$). We showed in section 2.3 the relationship between this parameter and the size of a spatial kernel: a larger parameter value induces smaller spatial neighbourhood influencing the measure and vice versa. In our approach, the scale-control property enables to obtain a similar regularisation level despite variations in brain size. To our opinion it is a crucial point which deserves more attention in neuroimaging where smoothing procedures are rarely used with an adaptive kernels for various brain sizes, except for instance kim_development_2016.
5.4 DPFDPF* does not compensate for variations beyond global scaling
Our approach is the first to explicitly allow for compensating for the influence of global brain size on local depth estimates. However, the increasing complexity of cortical folding during development lefevre_are_2016 or allometric variations with respect to brain size between adults toro_brain_2008; reardon_normative_2018 suggest that more subtle empirical laws exist between sulcal depth and the global size of the brain. Specifically, our observation of a strong influence of brain size (higher Wasserstein distance) on the DPFDPF* for the 200 smaller, younger subjects despite our formal normalization is expected. We interpret the sharp increase in variance for DPF* on this age range as the impact of the combination of increase in global size and magnitude of folding, that cannot be compensated by a global normalization factor. Regarding SULC, our experience 1 confirms that an increase in global size is sufficient to induce a spread in the distribution of the estimated depth, which is expected from a method that is not normalized. The observation of the increase in variance with age for very young babies is thus expected.
Normalizing for global scaling should thus not fully compensate for variations in sulcal depth. In this work, we propose an original formalization of the problem that is a contribution to addressing this question, but not in terms of biological processes. Additional studies are needed to better understand the interactions between biological processes and their dynamics, in order to approach the causal relationship between mechanisms acting at the cellular level, and the observed variations at the millimeter to centimeter scale (folds) and finally at the global brain scale. The ultimate sulcal depth estimation technique would eventually consider local normalization factors. Designing such a method allows us to also compensate for local variations in cortical geometry, for instance using the approach introduced in joshi_parameterization-based_2009. However, note that modelling variations in local cortical folding would amount to propose a model of brain folding like for instance tallinen_growth_2016. The inclusion of local normalization factors thus corresponds to a major change in the objectives and perspectives of the approach.
@@ 6 Conclusion
Ethics statement
This research study was conducted retrospectively using human subject data made available in open access by the dHCP (http://www.developingconnectome.org), the HCP (https://www.humanconnectome.org/) and the KKI (Kennedy Krieger Institute, landman_multi-parametric_2011). Each study was approved by the respective institutional review board. All procedures were performed in compliance with relevant laws and institutional guidelines and have been approved by the appropriate institutional committees. Informed consent was obtained for experimentation with human subjects.
@@ Acknowledgments
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used MISTRAL AI in the writing process to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ 6.6 Qualitative comparison on a spatio-temporal atlas
We provide on Fig.9 the results from Experiment#4 for the right hemisphere. The observations and interpretations are similar to these reported for the left hemisphere in section 4.4
Figure 9: Results from Experiment #4 for the right hemisphere. Top: Labeled sulcal regions shown on three time steps of the spatio-temporal atlas. The nomenclature of sulci and associated colors are shown on the right. The colors are indicative of the expected timing of formation: sulci colored in levels of blue are expected to appear before the sulci in red that appear before the sulci in green. Bottom: Curves showing the average sulcal depth estimated in the different sulcal regions for SULC (left) and $DPF_{abs}^{*}$ (Right).
In order to complement the results and observations from the Experiment#4, we show in Fig. 10 the sulcal depth maps obtained using $DPF^{*}$ and SULC on the surfacic atlas from the dHCP bozek_construction_2018 at 28, 32 and 44 weeks post conception. We observe a better dissociation between sulci and gyri at every age for $DPF^{*}$ compared to SULC. In addition, the estimated depth values are more homogeneous in gyral crests distributed across cortical regions with $DPF^{*}$, while with SULC we observe that high depth values are concentrated around large convex regions such as the anterior temporal lobe, the frontal pole and the occipital region. The $DPF^{*}$ gives a finer description of the folds, with a better definition of the small sulci located in the inferior temporal region, as well as in the medial side of the cortex (see for instance the small folds above the cingular sulcus).
Figure 10: Application of $DPF^{*}$ (left) and SULC (right) to the spatio-temporal atlas from bozek_construction_2018, computed from the dHCP dataset corresponding to 28, 32 and 44 weeks post-conception from top to bottom.
6.7 Influence of the spatial resolution of the mesh
The impact of the spatial resolution on the sulcal depth value estimated with $DPF^{*}$ is limited. To complement the quantitative assessment from Table 1, we provide on Fig.11 a visualization of the sulcal depth maps obtained after reducing the number of vertices to 75%, 50%, 25% and 10% for two subjects with different levels of folding.
Figure 11: Illustration of the $DPF^{*}$ computed for two individual surfaces after down sampling the number of vertices to 75%, 50%, and 25% of the original mesh.